A current of 3.40 A is passed through a Sn(NO3)2 solution. How long (in hours) would this current have to be applied to plate out 6.10 g of tin?
Answer #1

Sn(NO3)2 ---- > Sn2+ + 2NO3-

Sn2+ + 2e-   ---- > Sn

number of moles of Sn = 6.10 g / 118.71 g.mol^-1 = 0.05139 mole

1 mole of Sn requires 2 mole of electron so

0.05139 mole of Sn will require

= 0.05139 mole of Sn*( 2 mole of electron /1 mole of Sn)

= 0.1028 mole of electron

1 mole of electron carries 96500 C charge so

0.1028 mole of electron will carry 9920.2 C charge

C = A * sec

sec = C/A

sec = 9920.2/3.40 = 2917.7 sec

3600 sec = 1h so

2917.7 sec = 0.8105 h

Therefore, the time required is 0.8105 hours